OpenStudy (anonymous):

use the quadratic formula to find any x-intercepts on the graph of the equation. y=x^2-2x-2

OpenStudy (anonymous):

\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=1,b=-2,c=-2\]

OpenStudy (anonymous):

in fact for this one it is easier to "complete the square" \[x^2-2x-2=0\] \[x^2-2x=2\] \[(x-1)^2=2+1=3\] \[x-1=\pm\sqrt3\] \[x=1\pm\sqrt3\]